Given the cell notation, what is [Mn2+] if Ecell reaches 0.055 V? Mn(s) | Mn2+(aq) (0.090 M) || Cd2+(aq) (0.060 M) | Cd(s) at 30 °C

1
Expert's answer
2022-01-12T13:40:02-0500

"Cd_{(s)}\to\>Cd^{2+}_{(aq)}+2e^-\>\>\>E^{\theta}=0.40V"


"Mn^{2+}_{(aq)}+2e^-\to\>Mn_{(s)}\>\>\>E^{\theta}=-1.18V"


Total "=1.18-0.40=0.78V"


"E_{cell}=E^0-\frac{(0.0592)}{n}log\>Q"


Let "[Cd^{2+}]" decreases by "x"


"\implies[Mn^{2+}]" increases by "x"


"0.055=0.78-\frac{0.0592}{2}\>log\>\frac{0.090+x}{0.060-x}"


Solving


Log "\frac{0.090+x}{0.060-x}=24.493"


"\frac{0.090+x}{0.060-x}=3.112×10^{24}"


Ignoring "x" and "0.090" since they are negligible


"1.8672\u00d710^{23}=3.112\u00d710^{24}(x)"


"x=0.06"


"[Mn^{2+}]=0.06+0.09"

"=0.150M"